REGISTRATION FORM

PARTICIPANT

Family name First Name Institution Mailing address

Postal code  City  Country
Phone  Fax  e-mail

ACCOMPANYING PERSON(S)

Family name First Name

I intend to present a poster YES NO

I am enclosing my application form for a congress grant YES NO

REGISTRATION FEE

prior to March 15 March 16 — May 15 after May 16 and on site

Members of the EMS Non-members Accompanying persons

WAY OF PAYMENT Registration fee, one night hotel deposit and full payment of the tours should be paid by certified cheque or credit card. Each remittance must show the name of the participant and the following programme number: 4006–06–21076.

1. Cheque should be sent to: Malév Air Tours, Congress Department Roosevelt tér 2, H-1051 Budapest, Hungary

2. Credit card number expiry date

Diners Club  Visa  Master Card/Eurocard  American Expres

I authorise Malév Air Tours to charge my account with the registration fee, hotel deposit and optional programmes.

Date Signature

HOTEL RESERVATION

Choice  Name of hotel  Number of rooms  Date of
single  double  check in  lcheck out
1st  l
2nd  l
3rd  l

Hotel reservation will be made only upon receipt of hotel deposit which is one night’s fee of each booked room of your choice. OPTIONAL PROGRAMMES

Name of the tour  Date of the tour  Number of persons